OpenStudy (anonymous):

ABCD is a quadrilateral inscribed in a circle:

OpenStudy (anonymous):

What equation can be used to solve for angle B?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@agent0smith @AravindG @whpalmer4

OpenStudy (whpalmer4):

what do you know about the sum of the interior angles of a quadrilateral?

OpenStudy (whpalmer4):

hint: a square is a quadrilateral, too

OpenStudy (anonymous):

I know that The sum of any shape's interior angles (The angles inside the shape made where the sides meet) can be found using the equation (n-2)*180 where n=the number of sides the shape has. A quadrilateral has four sides so (4-2)*180 2*180 360 The sum of a quadrilateral's interior angles is 360 degrees.

OpenStudy (whpalmer4):

right. so you can write an equation for the sum of the angles and set it equal to 360, then solve for the value of x.

OpenStudy (anonymous):

these are the options by the way

OpenStudy (anonymous):

(x + 16) - (6x − 4) = 180 (x + 16) + (x) = 90 (2x + 16) + (x) = 180 (x + 16) − (x) = 90

OpenStudy (anonymous):

but none of these have 360 in them?

OpenStudy (anonymous):

im thinking it will be 2x + 16) + (x) = 180

OpenStudy (whpalmer4):

besides the interior angles summing to 360, the opposite angles will sum to 180

OpenStudy (anonymous):

yeah I checked them and the answer choices where right I coppied and pasted them

OpenStudy (whpalmer4):

yeah, I'd go with the 3rd one.

OpenStudy (anonymous):

nevermind Im just going to guess on this one can you help me with another question please

OpenStudy (whpalmer4):

I'm a little puzzled as to why the solution to that equation isn't giving us an answer that satisfies the sum of all 4 = 360, but maybe it's just a lack of coffee so far

OpenStudy (anonymous):

yeah lol its confusinf for me too

OpenStudy (whpalmer4):

well, what's the other question?